OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −(x + 6) y − 4 = 2(x + 6) y + 6 = −(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Slope: 2-(-6)/0-4 =8/-4 =2 equation: y-(-6)/x-4 =2 y+6/x-4 =2 y+6 = 2(x-4) I think it should be like this......

OpenStudy (anonymous):

Nah it was A

OpenStudy (anonymous):

Thanks for trying though:D

hartnn (hartnn):

slope =(y2-y1)/(x2-x21) = (0-4)/(2-(-6))

OpenStudy (anonymous):

OH! thanks for telling me the correct answer :-)

OpenStudy (anonymous):

Hahah no problem

hartnn (hartnn):

none of the choices are correct :\

hartnn (hartnn):

slope =(y2-y1)/(x2-x1) = (0-4)/(2-(-6)) = -1/2=m y-y1=m(x-x1) y-4=(-1/2)(x+6) or 2(y-4)=-(x+6)

OpenStudy (anonymous):

a is correct I know that for a fact buddy

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

None of the choices is correct. The right answer is y-4=(-1/2)(x+6). @hartnn is right.

OpenStudy (anonymous):

NONE ANS

OpenStudy (anonymous):

It's A, I took the test an got it right

hartnn (hartnn):

i believe in myself more than the software used in your test...i'd still say its not A , its 2(y-4)=-(x+6)

OpenStudy (anonymous):

well...

hartnn (hartnn):

and i have proved it to you....

OpenStudy (anonymous):

if you say so bro